zarkam21:

would this be 24 or 3x+17

Shadow:

Need a bit more information

zarkam21:

1 attachment
zarkam21:

Patience Batman :D

Vocaloid:

soo for this one you want to evaluate h(3) first, and then take the result of h(3) and plug that into g(x)

zarkam21:

zarkam21:

Here is the work I did just want to know if I am right and if not what I did wrong,, really want to understand the concept

Vocaloid:

first let's just look at h for now

zarkam21:

okay

Vocaloid:

if h(x) = sqrt(x) and x = 9, what is h?

zarkam21:

3

Vocaloid:

good, so this is the result of h(9) and we plug in "3" into g(x)

Vocaloid:

g(x) = 2x - 1 = what when x = 3?

zarkam21:

so g(3) right

Vocaloid:

yes, and this equals?

zarkam21:

g(3)=6-1

zarkam21:

5?

Vocaloid:

good, that's the final answer

zarkam21:

Oh that was easy

zarkam21:

Thank you I'm gonna write this down for future reference

zarkam21:

Quick question

zarkam21:

Doesn't it say find goh though,, or is this finding that

Vocaloid:

yeah, this is the same as finding g dot h

zarkam21:

Okay